# Feasibility-Aware Decision-Focused Learning for Predicting Parameters in the Constraints

### v1

@@ 4.1 Loss Function to Penalize for Infeasibility-Inducing Predictions
Note that $\mathit{UNSAT}_{i}\left(\bm{x}_{\mathit{neg}},\bm{\rho}_{\kappa}\right)$ can be viewed as a mask that equals to $1$ when $\bm{x}_{\mathit{neg}}$ violates constraint $i$ under the true parameter $\bm{\rho}_{\kappa}$. Using the definition of $\mathit{P}_{\theta}\left(\mathit{UNSAT}_{i}\left(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\middle|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)$ from Eq. 6, we can write the following:
$\log\mathit{P}_{\theta}\Big(\mathit{UNSAT}_{i}\left(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\Big|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\Big)=-\log\left(1+\exp\left(-g_{i}\left(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}\left(\bm{\phi}_{\kappa}\right)\right)\right)\right)$
Next we will propose a loss function by minimizing the negative log-likelihood. Furthermore, to encourage the constraints to be violated with a buffer, we propose adding a margin parameter $\Upsilon>0$. Since this loss penalizes cases where the $\bm{x}_{\mathit{neg}}$ solution is infeasible, we refer to it as Infeasibility Aware Loss (IAL).
@@
where we write $\log 1+\exp(.)$ as $\mathit{softplus}(.)$.
@@
From this expression, we see that for a truly violated constraint, i.e., of which $\mathit{UNSAT}_{i}\big(\bm{x}_{\mathit{neg}},\bm{\rho}_{\kappa}\big)=1$, the loss becomes zero only if $g_{i}\big(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\big)>\Upsilon$. In other words, the predicted constraint must also consider this constraint to be violated. $\overline{\mathcal{L}^{\mathit{IAL}}}(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa}))$ is zero, if this holds for all constraints violated by $\bm{x}_{\mathit{neg}}$.
@@
The IAL loss formulation is based on $\bm{x}_{\mathit{neg}}$, a candidate solution which violates at least one constraint with respect to the true parameter $\bm{\rho}_{\kappa}$. This leads to the question of how to obtain such a $\bm{x}_{\mathit{neg}}$. We adopt the following approach: for each instance, after predicting $\bm{\hat{\rho}}_{\kappa}$, we solve the COP using $\bm{\hat{\rho}}_{\kappa}$ and obtain a solution, $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa})$. If this solution is infeasible under the true parameters, $\bm{\rho}_{\kappa}$, we use it as $\bm{x}_{\mathit{neg}}$. Otherwise, we set the loss to zero, since the predicted solution satisfies all the true constraints, which is the preferred outcome. We call this final loss function infeasibility preserving loss (IPL), $\mathcal{L}^{\mathit{IPL}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$.
@@
For a true parameter $\bm{\rho}_{\kappa}$ and a predicted parameter, $\bm{\hat{\rho}}_{\kappa}$, $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ is zero if and only if $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$, a solution of $\bm{\hat{\rho}}_{\kappa}$, is feasible with respect to $\bm{\rho}_{\kappa}$.
@@ 4.2 Loss Function for Preserving Optimal Solutions
The above loss penalizes predicted parameters when the predicted solution becomes infeasible with respect to the true parameter. Minimizing only $\mathcal{L}^{\mathit{IPL}}$ can lead to learning a feasible region that is much stricter than the true feasible region, as any solution therein will be truly feasible. This might lead to learning too strict parameters, not practical. For example, when learning the capacity parameter in a knapsack problem, minimizing only $\mathcal{L}^{\mathit{IPL}}$ might result in capacity values so small that no items are selected. While such solutions are always feasible with respect to the true parameters, they are suboptimal and impractical. Therefore, we also need a loss function to penalize if the learned parameter turns out to be too strict. To do this, we define a loss function that penalizes predicted parameters when they render known optimal solutions infeasible. Within the MLE framework, this corresponds to maximizing the likelihood that the true solution is feasible.
Thus, we formulate the probability of an assignment, $\bm{x}$ being a feasible point, given the parameter $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$. Note that $\bm{x}$ can be a feasible assignment, only if it satisfies all the constraints. While the probabilities of satisfying constraints $i$ and $j$ are not independent, they are conditionally independent [5, Chapter 8] given the constraint parameter vector $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$. This is because the constraint parameter vector determines the feasibility of each constraint independently. Once $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$ is known, the probability of satisfying constraint $i$ depends only on $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$, and any additional information about constraint $j$ being satisfied does not affect that probability. Because of the conditional independence, we can express the conditional probability of the assignment, $\bm{x}$ being feasible given $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$, by:
@@
We point out that in the training data, we know that $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\rho}_{\kappa})$ is a feasible solution. Denoting $\bm{x}^{\star}_{\kappa}$ as the known feasible point, we can formulate a loss function by minimizing the negative log-likelihood. In our implementation, we add a margin parameter, $\Upsilon$ as we have done for $\mathcal{L}^{\mathit{IPL}}$. Since this loss is designed to maximize the probability that the optimal solution remains feasible, we refer to it as the Optimality-Preserving Loss (OPL). We express it in the following form:
@@
where we write $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$ as $\bm{\hat{\rho}}_{\kappa}$ for brevity. From this expression, it is easy to see that $\mathcal{L}^{\mathit{OPL}}$ attains its minimum value of zero when all constraint values less than or equal to $\Upsilon$, i.e., when all the constraints are satisfied with some margin.
@@
For a true parameter $\bm{\rho}_{\kappa}$ and a predicted parameter, both $\overline{\mathcal{L}^{\mathit{OPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$, and $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ will be zero if and only if, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$, a solution of $\bm{\hat{\rho}}_{\kappa}$, is an optimal solution with respect to $\bm{\rho}_{\kappa}$.
@@
Let $\overline{\mathcal{L}^{\mathit{OPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ and $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ are equal to zero. We have proved in Lemma 1, that $\overline{\mathcal{L}^{\mathit{IPL}}}$ being zero implies that a solution of $\bm{\hat{\rho}}_{\kappa}$ is feasible with respect to $\kappa$. So, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$, a solution of $\bm{\hat{\rho}}_{\kappa}$ lies in the set $\mathcal{F}(\bm{\rho}_{\kappa})$.
@@
As, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)\in\mathcal{F}(\bm{\rho}_{\kappa})$, from the above argument, we can write
@@
As, $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ is a feasible solution with respect to $\bm{\rho}_{\kappa}$ and its objective value is equal to the optimal objective value, it is an optimal solution.
Now, to prove in the reverse direction, assume that $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is an optimal solution with respect to $\bm{\rho}_{\kappa}$. This definitely implies that $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is a feasible solution with respect to $\bm{\rho}_{\kappa}$. Therefore, from Lemma 1, $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})=0$. As, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ adheres to all the constraints when $\bm{\hat{\rho}}_{\kappa}$ is the constraints parameter value,
@@
As, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is an optimal solution with respect to $\bm{\rho}_{\kappa}$, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is one of the $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\rho}_{\kappa}\right)$. If we replace $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ with $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\rho}_{\kappa}\right)$ in Eq. 20, $\overline{\mathcal{L}^{\mathit{OPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})=0$ by definition ( Eq. 17). ∎
@@ 4.3 Minimizing a Convex Combination of $\mathcal{L}^{\mathit{OPL}}$ and $\mathcal{L}^{\mathit{IPL}}$
Lemma 2 states that $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ is an optimal solution with respect to $\bm{\rho}_{\kappa}$, if both $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$ are zero. We point out that $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$ represent two conflicting goals. Minimizing only $\mathcal{L}^{\mathit{OPL}}$ encourages the model to predict loose constraints so that all true feasible solutions remain feasible, hence making even infeasible solutions appear feasible. On the other hand, minimizing $\mathcal{L}^{\mathit{IPL}}$ encourages predicting strict constraints. If the learned parameters are too strict, even the true optimal points may become infeasible. We propose a simple straightforward to minimize a convex combination of them, in the following form:
@@
$\alpha\in[0,1]$ is a hyperparameter we call the infeasibility-aversion coefficient. It reflects how much importance the decision-maker places on avoiding infeasible solutions (through IPL) versus maintaining the feasibility of known feasible ones (through OPL). The two extremes are $\alpha=1$ and $\alpha=0$, both of which represent unrealistic preferences. An $\alpha=1$ means the decision-maker prioritizes only feasibility, possibly ending up with trivial solutions (e.g., picking no items in a knapsack). An $\alpha$ near 0 implies a focus on high-reward solutions, even if they are infeasible (e.g., selecting all items).
In this section, we have proposed a novel approach for ‘optimizing decision through end-to-end constraint estimation’ (Odece), involving two novel loss functions – $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$. We have proposed to consider a weighted average of the two losses to reflect the decision-maker’s subjective preference of infeasibility over siboptimality. Next, we will experimentally investigate how Odece performs compared to other existing approaches for predicting constraint parameters.

### v2

@@ 4.1 Loss Function to Penalize Infeasibility-Inducing Predictions
Note that $\mathit{UNSAT}_{i}\left(\bm{x}_{\mathit{neg}},\bm{\rho}_{\kappa}\right)$ can be viewed as a mask that equals $1$ when $\bm{x}_{\mathit{neg}}$ violates constraint $i$ under the true parameters $\bm{\rho}_{\kappa}$. Using the definition of $\mathit{P}_{\theta}\left(\mathit{UNSAT}_{i}\left(\bm{x},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\middle|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)$ from Eq. 6, we can write the following:
$\log\mathit{P}_{\theta}\Big(\mathit{UNSAT}_{i}\left(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\right)\Big|\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\Big)=-\log\left(1+\exp\left(-g_{i}\left(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}\left(\bm{\phi}_{\kappa}\right)\right)\right)\right)$ (10)
We now propose a loss function to maximize this likelihood (i.e., minimize the negative log-likelihood). We include a margin parameter $\Upsilon>0$ to encourage the constraints to be violated with a buffer. Since this loss penalizes cases where an infeasible assignment $\bm{x}_{\mathit{neg}}$ is considered feasible under the predicted parameters, we refer to it as the Infeasibility-Aware Loss (IAL).
@@
where we write $\log(1+\exp(.))$ as $\mathit{softplus}(.)$.
@@
From this expression, we see that for a truly violated constraint, i.e., of which $\mathit{UNSAT}_{i}\big(\bm{x}_{\mathit{neg}},\bm{\rho}_{\kappa}\big)=1$, the loss becomes zero only if $g_{i}\big(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})\big)>\Upsilon$. In other words, the predicted constraint must also be violated. $\overline{\mathcal{L}^{\mathit{IAL}}}(\bm{x}_{\mathit{neg}},\mathcal{M}_{\theta}(\bm{\phi}_{\kappa}))$ is zero if this holds for all constraints violated by $\bm{x}_{\mathit{neg}}$.
@@
The IAL loss formulation is based on $\bm{x}_{\mathit{neg}}$, a candidate assignment which violates at least one constraint with respect to the true parameter $\bm{\rho}_{\kappa}$. This leads to the question of how to obtain such a $\bm{x}_{\mathit{neg}}$. We adopt the following approach: for each instance, after predicting $\bm{\hat{\rho}}_{\kappa}$, we solve the COP using $\bm{\hat{\rho}}_{\kappa}$ and obtain a solution $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa})$. If this solution is infeasible under the true parameters $\bm{\rho}_{\kappa}$, we use it as $\bm{x}_{\mathit{neg}}$. Otherwise, we set the loss to zero, since the predicted solution satisfies all the true constraints, which is the preferred outcome. We call this final loss function the Infeasibility Penalty Loss (IPL), as it penalizes the predicted parameters when the resulting predicted solution violates the true constraints.
@@
For a true parameter $\bm{\rho}_{\kappa}$ and a predicted parameter $\bm{\hat{\rho}}_{\kappa}$, $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ is zero if and only if $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$, an optimal solution for $\bm{\hat{\rho}}_{\kappa}$, is feasible with respect to $\bm{\rho}_{\kappa}$.
@@ 4.2 Loss Function for Preserving Optimal Solutions
The IPL loss penalizes predicted parameters when the predicted solution becomes infeasible with respect to the true parameters. However, minimizing only $\mathcal{L}^{\mathit{IPL}}$ can lead to learning a feasible region that is much stricter than the true feasible region, as any solution therein will be feasible under the true parameters. For example, when learning the capacity parameter in a knapsack problem, minimizing only $\mathcal{L}^{\mathit{IPL}}$ might result in capacity values so small that no items are selected. While such solutions are always feasible with respect to the true parameters, they are suboptimal and of no practical value. Therefore, we also need a loss function to penalize cases where the learned parameters turn out to be too strict. To do this, we define a loss function that penalizes predicted parameters when they render known optimal solutions infeasible. Within the MLE framework, this corresponds to maximizing the likelihood that the true solution is feasible.
Thus, we formulate the probability of an assignment $\bm{x}$ being a feasible point, given the parameter $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$. Note that $\bm{x}$ is a feasible assignment only if it satisfies all the constraints. While the probabilities of satisfying constraints $i$ and $j$ are not independent, they are conditionally independent [7, Chapter 8] given the constraint parameter vector $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$. This is because the constraint parameter vector determines the feasibility of each constraint independently. Once $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$ is known, the probability of satisfying constraint $i$ depends only on $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$, and any additional information about constraint $j$ being satisfied does not affect that probability. Because of the conditional independence, we can express the conditional probability of the assignment, $\bm{x}$ being feasible given $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$, by:
@@
We point out that in the training data, we know that $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\rho}_{\kappa})$ is a feasible solution. Denoting $\bm{x}^{\star}_{\kappa}$ as the known feasible point, we can formulate a loss function by minimizing the negative log-likelihood. In our implementation, we add a margin parameter $\Upsilon$ as we have done for $\mathcal{L}^{\mathit{IPL}}$. Since this loss is designed to maximize the probability that the optimal solution remains feasible, we refer to it as the Optimality-Preserving Loss (OPL). We express it in the following form:
@@
where we write $\mathcal{M}_{\theta}(\bm{\phi}_{\kappa})$ as $\bm{\hat{\rho}}_{\kappa}$ for brevity. From this expression, it is easy to see that $\mathcal{L}^{\mathit{OPL}}$ attains its minimum value of zero when all constraint values are less than or equal to $\Upsilon$, i.e., when all the constraints are satisfied with some margin.
@@
For a true parameter $\bm{\rho}_{\kappa}$ and a predicted parameter $\bm{\hat{\rho}}_{\kappa}$, both $\overline{\mathcal{L}^{\mathit{OPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ and $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ will be zero if and only if $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$, a solution of $\bm{\hat{\rho}}_{\kappa}$, is an optimal solution with respect to $\bm{\rho}_{\kappa}$.
@@
Let $\overline{\mathcal{L}^{\mathit{OPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ and $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ be equal to zero. We have proved in Lemma 1 that $\overline{\mathcal{L}^{\mathit{IPL}}}$ being zero implies that a solution of $\bm{\hat{\rho}}_{\kappa}$ is feasible with respect to $\kappa$. So, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$, a solution of $\bm{\hat{\rho}}_{\kappa}$ lies in the set $\mathcal{F}(\bm{\rho}_{\kappa})$.
@@
As $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)\in\mathcal{F}(\bm{\rho}_{\kappa})$, from the above argument, we can write
@@
As $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa})$ is a feasible solution with respect to $\bm{\rho}_{\kappa}$ and its objective value is equal to the optimal objective value, it is an optimal solution.
Now, to prove the opposite direction, assume that $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is an optimal solution with respect to $\bm{\rho}_{\kappa}$. This implies that $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is a feasible solution with respect to $\bm{\rho}_{\kappa}$. Therefore, from Lemma 1, $\overline{\mathcal{L}^{\mathit{IPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})=0$. As $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ adheres to all the constraints when $\bm{\hat{\rho}}_{\kappa}$ is the constraints parameter value,
@@
As $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is an optimal solution with respect to $\bm{\rho}_{\kappa}$, $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ is one of the $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\rho}_{\kappa}\right)$. If we replace $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa}\right)$ with $\bm{x}^{\star}\left(\bm{q}_{\kappa},\bm{\rho}_{\kappa}\right)$ in Eq. 21, $\overline{\mathcal{L}^{\mathit{OPL}}}(\bm{\rho}_{\kappa},\bm{\hat{\rho}}_{\kappa})=0$ by definition (Eq. 18). ∎
@@ 4.3 Minimizing a Convex Combination of $\mathcal{L}^{\mathit{OPL}}$ and $\mathcal{L}^{\mathit{IPL}}$
Lemma 2 suggests that by minimizing both $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$ to zero, we can obtain a solution $\bm{x}^{\star}(\bm{q}_{\kappa},\bm{\hat{\rho}}_{\kappa})$, which is optimal with respect to $\bm{\rho}_{\kappa}$. However, simultaneously minimizing $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$ presents a challenge. This is because $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$ represent two conflicting goals. Minimizing only $\mathcal{L}^{\mathit{OPL}}$ encourages the model to predict loose constraints so that all true feasible solutions remain feasible, hence making even infeasible solutions appear feasible. Conversely, minimizing $\mathcal{L}^{\mathit{IPL}}$ encourages stricter constraints, which may render even true optimal solutions infeasible. We thus propose to minimize a convex combination of these losses, in the following form:
@@
$\alpha\in[0,1]$ is a hyperparameter we call the infeasibility-aversion coefficient. It reflects how much importance the decision-maker places on avoiding infeasible solutions (through IPL) versus maintaining the feasibility of known feasible ones (through OPL). The two extremes are $\alpha=1$ and $\alpha=0$, both of which represent unrealistic preferences. $\alpha=1$ means the decision-maker prioritizes only feasibility, possibly ending up with trivial solutions (e.g., picking no items in a knapsack problem). $\alpha$ near 0 implies a focus on high-reward solutions, even if they are infeasible (e.g., selecting all items).
In this section, we have proposed a novel approach for ‘optimizing decisions through end-to-end constraint estimation’ (Odece), involving two novel loss functions: $\mathcal{L}^{\mathit{IPL}}$ and $\mathcal{L}^{\mathit{OPL}}$. We have proposed to consider a weighted average of the two losses to reflect the decision-maker’s subjective preference of infeasibility over suboptimality. Next, we will experimentally investigate how Odece performs compared to other existing approaches for predicting constraint parameters.